# Why Policy Gradient Algorithms Work for Undiscounted Total-Reward MDPs

### v1

@@ Appendix C Omitted proofs in Section 6
Following the proof strategy of Xiao (2022), we consider mirror descent framework.
@@
Following the derivations of Shani et al. (2020); Xiao (2022), we consider policy mirror descent methods with dynamically weighted divergences:
$\pi_{k+1}=\operatorname*{arg\,min}_{\pi\in\Pi}\left\{-\eta_{k}\big\langle\nabla V_{\mu}^{\pi_{k}},\,\pi\big\rangle+\,D_{\delta_{\mu}(\pi_{k})}\!\big(\pi,\pi_{k}\big)\right\},$
@@
Consider direction parametrization, and by Theorem 1, we have
$\pi_{k+1}$ $=\operatorname*{arg\,min}_{\pi\in\Pi}\left\{-\eta_{k}\sum_{s\in\mathcal{S}}\delta^{\pi_{k}}_{\mu}(s)\big(\sum_{s\in\mathcal{S}}Q^{\pi_{k}}(s,a)\pi(a\,|\,s)+D\big(\pi(\cdot\,|\,s),\pi_{k}(\cdot\,|\,s)\big)\right\}$
$=\operatorname*{arg\,min}_{\pi\in\Pi}\left\{-\eta_{k}\sum_{s\in\mathcal{S}}\big(\sum_{s\in\mathcal{S}}Q^{\pi_{k}}(s,a)\pi(a\,|\,s)+D\big(\pi(\cdot\,|\,s),\pi_{k}(\cdot\,|\,s)\big)\right\}$
@@
$\pi_{k+1}(\cdot\,|\,s)=\operatorname*{arg\,min}_{p\in\mathcal{M}(\mathcal{A})}\Big\{-\eta_{k}\ \sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)p(a)+D\big(p,\,\pi_{k}(\cdot\,|\,s)\big)\Big\}$
@@
C.1 Proof of Lemma 7
We proved more detailed version of Lemma 7.
Lemma 11.
Under Assumption 1, for a given $\mu$ with full support, the natural policy gradient with step size $\eta_{k}>0$ generates a sequence of policies $\{\pi_{k}\}^{\infty}_{k=1}$ satisfying
@@ C.1 Proof of Lemma 7
Proof of Lemma 7.
Applying Fact 4 with $\mathcal{C}=\mathcal{M}(\mathcal{A})$, $\phi(p)=-\eta_{k}\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)p(a)$, and KL divergence as Bregman divergence, we obtain
$\eta_{k}\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)p(a)+\text{KL}\big(\pi_{k+1}(\cdot\,|\,s),\,\pi_{k}(\cdot\,|\,s)\big)$
$\leq\eta_{k}\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)\pi_{k+1}(a\,|\,s)+\text{KL}\big(p,\pi_{k}(\cdot\,|\,s)\big)-\text{KL}\big(p,\pi_{k+1}(\cdot\,|\,s)\big)$
@@
$\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)(p(a)-\pi_{k+1}(a\,|\,s))+\frac{1}{\eta_{k}}\text{KL}\!\big(\pi_{k+1}(\cdot\,|\,s),\pi_{k}(\cdot\,|\,s)\big)$
@@
$\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))\leq-\frac{1}{\eta_{k}}\text{KL}\big(\pi_{k+1}(\cdot\,|\,s),\pi_{k}(\cdot\,|\,s)\big)-\frac{1}{\eta_{k}}\text{KL}\big(\pi_{k}(\cdot\,|\,s),\pi_{k+1}(\cdot\,|\,s)\big).$
Then, first results comes from nonnegativity of Bregman divergence and second result comes from Corollary 5. ∎
C.2 Proof of Theorem 3
@@
$\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))+\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)(\pi(a\,|\,s)-\pi_{k}(a\,|\,s))$
@@
$\sum_{s\in\mathcal{S}}\frac{\delta_{\mu}^{\pi}(s)}{\left\|{\delta_{\mu}^{\pi}}\right\|_{\infty}}\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))$
$\geq\sum_{s\in\mathcal{S}}(\delta_{\mu}^{\pi})^{\prime}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))$
$\geq\sum_{s\in\mathcal{S}}\delta_{(\delta_{\mu}^{\pi})^{\prime}}^{\pi_{k+1}}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))$
@@
where $(\delta_{\mu}^{\pi})^{\prime}$ is probability distribution satisfying $(\delta_{\mu}^{\pi})^{\prime}(s)\geq\frac{\delta_{\mu}^{\pi}(s)}{\left\|{\delta_{\mu}^{\pi}}\right\|_{\infty}}$, the first and second inequalities come from Lemma 11, and the last equality comes from Corollary 5.
For the second term, by Corollary 5,
$\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)(\pi(a\,|\,s)-\pi_{k}(a\,|\,s))=V_{\mu}^{\pi}-V_{\mu}^{\pi_{k}}.$
@@
Since $V_{\mu}^{\pi_{k}}$ is non-decreasing by Lemma 7 and KL-divergence is non-negative, we conclude that
@@
C.3 Proof of Corollary 2
@@
In the previous proof of Theorem 3, let $\pi=\pi^{\star}$, and use Corollary 6 instead of Corollary 5 and the fact that $V^{\pi}_{\mu}\geq 0$. Then, we obtain desired result. ∎
C.4 Proof of Theorem 4
@@
Lemma 12.
Under Assumption 1, for a given $\mu$ with full support, the natural policy gradient with step size $\eta_{k}>0$ generates a sequence of policies $\{\pi_{k}\}^{\infty}_{k=1}$ satisfying,
@@
Proof of Lemma 12.
In the previous proof of Theorem 3, we showed that
$\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))+\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)(\pi(a\,|\,s)-\pi_{k}(a\,|\,s))$
@@
$\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))$
$=\sum_{s\in\mathcal{S}}\frac{\delta_{\mu}^{\pi}}{\delta_{\mu}^{\pi_{k+1}}}\,\delta_{\mu}^{\pi_{k+1}}\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))$
$\geq\left\lVert\frac{\delta_{\mu}^{\pi}}{\delta_{\mu}^{\pi_{k+1}}}\right\rVert_{\infty}\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi_{k+1}}\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))$
@@
where the first inequality is from Lemma 7, and the last equality comes from Corollary 5.
For the second term, by Corollary 5,
$\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)(\pi(a\,|\,s)-\pi_{k}(a\,|\,s))=V_{\mu}^{\pi}-V_{\mu}^{\pi_{k}}.$
We obtain desired result after substitution.
We are now ready to prove Theorem 4
@@
Since $\delta_{\mu}^{\pi_{k}}(s)\geq\mu(s)$ for all $s\in\mathcal{S}$, $\vartheta^{\pi}_{k}\leq\vartheta^{\pi}_{\mu}$ for all $k\geq 0$. $U^{\pi}_{k+1}-U^{\pi}_{k}\leq 0$ for all $k\geq 0$ by Lemma 11, and by Lemma 12,

### v2

@@ Appendix C Omitted proofs in Section 6
C.1 Reformulation of natural policy gradient
We basically follow the derivation in (Agarwal et al., 2021). First, we provide explicit form of policy gradient with softmax parametrization.
Lemma C.1.
$\frac{\partial V^{\pi_{\theta}}_{\mu}}{\partial\theta_{s,a}}=\delta_{\mu}^{\pi}(s)\pi_{\theta}(a\mid s)(Q^{\pi_{\theta}}(s,a)-V^{\pi_{\theta}}(s)).$
Proof.
With the softmax policy parameterization, we have
$\frac{\partial\log\pi_{\theta}(a\mid s)}{\partial\theta_{s^{\prime},a^{\prime}}}=\mathbf{1}[s=s^{\prime}]\left(\mathbf{1}[a=a^{\prime}]-\pi_{\theta}(a^{\prime}\mid s)\right)$
where $\mathbf{1}$ is the indicator function. Then, by Theorem 4.7,
$\frac{\partial V^{\pi_{\theta}}_{\mu}}{\partial\theta_{s^{\prime},a^{\prime}}}$ $=\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi}(s^{\prime})\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid s)}\biggl[Q^{\pi_{\theta}}(s,a)\,\frac{\partial\log\pi_{\theta}(a\mid s)}{\partial\theta_{s^{\prime},a^{\prime}}}\biggr]$
$=\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi}(s^{\prime})\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid s)}\bigl[Q^{\pi_{\theta}}(s,a)\,\mathbf{1}[s=s^{\prime}]\bigl(\mathbf{1}[a=a^{\prime}]-\pi_{\theta}(a^{\prime}\mid s)\bigr)\bigr]$
$=\delta_{\mu}^{\pi}(s^{\prime})\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid s^{\prime})}\bigl[Q^{\pi_{\theta}}(s^{\prime},a)\bigl(\mathbf{1}[a=a^{\prime}]-\pi_{\theta}(a^{\prime}\mid s^{\prime})\bigr)\bigr]$
$=\delta_{\mu}^{\pi}(s^{\prime})\Bigl(\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid s^{\prime})}\bigl[Q^{\pi_{\theta}}(s^{\prime},a)\mathbf{1}[a=a^{\prime}]\bigr]-\pi_{\theta}(a^{\prime}\mid s^{\prime})\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid s^{\prime})}\bigl[Q^{\pi_{\theta}}(s^{\prime},a)\bigr]\Bigr)$
$=\delta_{\mu}^{\pi}(s^{\prime})\,\pi_{\theta}(a^{\prime}\mid s^{\prime})(Q^{\pi_{\theta}}(s^{\prime},a^{\prime})-V^{\pi_{\theta}}(s^{\prime})).$
Now, we derive the reformulation of natural policy gradient.
Proof.
First, we have
$w^{\top}\nabla_{\theta}\log\pi_{\theta}(a\mid s)=w_{s,a}-\sum_{a^{\prime}\in\mathcal{A}}w_{s,a^{\prime}}\pi_{\theta}(a^{\prime}\mid s).$
Let $\overline{w}_{s}=\sum_{a^{\prime}\in\mathcal{A}}w_{s,a^{\prime}}\pi_{\theta}(a^{\prime}\mid s)$, which is independent of $a$. Then,
$\mathcal{F}_{\mu}(\theta)w$ $=\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi_{\theta}}(s)\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid s)}\bigl[\nabla_{\theta}\log\pi_{\theta}(a\mid s)\bigl(w^{\top}\nabla_{\theta}\log\pi_{\theta}(a\mid s)\bigr)\bigr]$
$=\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi_{\theta}}(s)\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid s)}\bigl[\nabla_{\theta}\log\pi_{\theta}(a\mid s)\bigl(w_{s,a}-\overline{w}_{s}\bigr)\bigr]$
$=\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi_{\theta}}(s)\mathbb{E}_{a\sim\pi_{\theta}(\cdot\mid s)}\bigl[w_{s,a}\,\nabla_{\theta}\log\pi_{\theta}(a\mid s)\bigr],$
where the last equality uses log derivative trick with $\overline{w}_{s}$, and this implies that
$\bigl[\mathcal{F}_{\mu}(\theta)w\bigr]_{s^{\prime},a^{\prime}}=\delta_{\mu}^{\pi_{\theta}}(s^{\prime})\,\pi_{\theta}(a^{\prime}\mid s^{\prime})\bigl(w_{s^{\prime},a^{\prime}}-\overline{w}_{s^{\prime}}\bigr).$
By property of Moore–Penrose pseudoinverse, $\bigl(\mathcal{F}_{\mu}(\theta)\bigr)^{\dagger}\nabla V^{\pi_{\theta}}_{\mu}$ is the minimum-norm solution of $\min_{w}\bigl\|\nabla V^{\pi_{\theta}}_{\mu}-\mathcal{F}_{\mu}(\theta)w\bigr\|_{2}^{2}$ where $\|\cdot\|_{2}$ is Euclidean norm. By lemma C.1, we have
$\bigl\|\nabla V^{\pi_{\theta}}_{\mu}-\mathcal{F}_{\mu}(\theta)w\bigr\|^{2}$ $=\sum_{s,a}\left(\delta_{\mu}^{\pi_{\theta}}(s)\pi_{\theta}(a\mid s)\left(Q^{\pi_{\theta}}(s,a)-V^{\pi_{\theta}}(s)-w_{s,a}-\sum_{a^{\prime}\in\mathcal{A}}w_{s,a^{\prime}}\pi_{\theta}(a^{\prime}\mid s)\right)\right)^{2}.$
If $w=Q^{\pi_{\theta}}(s,a)-V^{\pi_{\theta}}(s)$, $\bigl\|\nabla V^{\pi_{\theta}}_{\mu}-\mathcal{F}_{\mu}(\theta)w\bigr\|^{2}=0$, and $w$ has the form of $Q^{\pi_{\theta}}(s,a)+v(s)$ where $v$ only depends on state. Therefore, plugging $\mathcal{F}_{\mu}(\theta)^{\dagger}\nabla V^{\pi_{\theta}}_{\mu}=Q^{\pi_{\theta}}(s,a)+v(s)$ into the original form of natural policy gradient, we obtain reformulation of natural policy gradient.
Following the proof strategy of (Xiao, 2022), we consider mirror descent framework.
@@ C.1 Reformulation of natural policy gradient
Following the derivations of (Shani et al., 2020; Xiao, 2022), we consider policy mirror descent methods with dynamically weighted divergences:
$\pi_{k+1}=\operatorname{argmin}_{\pi\in\Pi}\left\{-\eta_{k}\big\langle\nabla V_{\mu}^{\pi_{k}},\,\pi\big\rangle+\,D_{\delta_{\mu}(\pi_{k})}\!\big(\pi,\pi_{k}\big)\right\},$
@@
Consider direction parametrization, and by Theorem 4.7, we have
$\pi_{k+1}$ $=\operatorname{argmin}_{\pi\in\Pi}\left\{-\eta_{k}\sum_{s\in\mathcal{S}}\delta^{\pi_{k}}_{\mu}(s)\big(\sum_{s\in\mathcal{S}}Q^{\pi_{k}}(s,a)\pi(a\,|\,s)+D\big(\pi(\cdot\,|\,s),\pi_{k}(\cdot\,|\,s)\big)\right\}$
$=\operatorname{argmin}_{\pi\in\Pi}\left\{-\eta_{k}\sum_{s\in\mathcal{S}}\big(\sum_{s\in\mathcal{S}}Q^{\pi_{k}}(s,a)\pi(a\,|\,s)+D\big(\pi(\cdot\,|\,s),\pi_{k}(\cdot\,|\,s)\big)\right\}$
@@
$\pi_{k+1}(\cdot\,|\,s)=\operatorname{argmin}_{p\in\mathcal{M}(\mathcal{A})}\Big\{-\eta_{k}\ \sum_{a\in\mathcal{A}}Q^{\pi_{k}}(a,s)p(a)+D\big(p,\,\pi_{k}(\cdot\,|\,s)\big)\Big\}$
@@
C.2 Proof of Lemma 6.1
We proved more detailed version of Lemma 6.1.
Lemma C.2.
Under Assumption 2.1, for given $\pi_{0}\in\Pi_{+}$ and $\mu$ with full support, the natural policy gradient with step size $\eta_{k}>0$ generates a sequence of policies $\{\pi_{k}\}^{\infty}_{k=1}$ satisfying
@@
Proof of Lemma 6.1.
Applying Fact 4 with $\mathcal{C}=\mathcal{M}(\mathcal{A})$, $\phi(p)=-\eta_{k}\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(s,a)p(a)$, and KL divergence as Bregman divergence, we obtain
$\eta_{k}\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(s,a)p(a)+\text{KL}\big(\pi_{k+1}(\cdot\,|\,s),\,\pi_{k}(\cdot\,|\,s)\big)$
$\leq\eta_{k}\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(s,a)\pi_{k+1}(a\,|\,s)+\text{KL}\big(p,\pi_{k}(\cdot\,|\,s)\big)-\text{KL}\big(p,\pi_{k+1}(\cdot\,|\,s)\big)$
@@
$\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(s,a)(p(a)-\pi_{k+1}(a\,|\,s))+\frac{1}{\eta_{k}}\text{KL}\!\big(\pi_{k+1}(\cdot\,|\,s),\pi_{k}(\cdot\,|\,s)\big)$
@@
$\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(s,a)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))\leq-\frac{1}{\eta_{k}}\text{KL}\big(\pi_{k+1}(\cdot\,|\,s),\pi_{k}(\cdot\,|\,s)\big)-\frac{1}{\eta_{k}}\text{KL}\big(\pi_{k}(\cdot\,|\,s),\pi_{k+1}(\cdot\,|\,s)\big).$
Then, the first results comes from nonnegativity of Bregman divergence and second result comes from Corollary A.2. ∎
C.3 Proof of Theorem 6.2
@@
$\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(s,a)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))+\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(s,a)(\pi(a\,|\,s)-\pi_{k}(a\,|\,s))$
@@
$\sum_{s\in\mathcal{S}}\frac{\delta_{\mu}^{\pi}(s)}{\|\delta_{\mu}^{\pi}\|_{1}}\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(s,a)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))$
$\geq\sum_{s\in\mathcal{S}}(\delta_{\mu}^{\pi})^{\prime}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(s,a)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))$
$\geq\sum_{s\in\mathcal{S}}\delta_{(\delta_{\mu}^{\pi})^{\prime}}^{\pi_{k+1}}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(s,a)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))$
@@
where $(\delta_{\mu}^{\pi})^{\prime}$ is probability distribution satisfying $(\delta_{\mu}^{\pi})^{\prime}(s)=\frac{\delta_{\mu}^{\pi}(s)}{\left\|{\delta_{\mu}^{\pi}}\right\|_{\infty}}$, the first and second inequalities come from Lemma C.2, and the last equality comes from Corollary A.2.
For the second term, by Corollary A.2,
$\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(s,a)(\pi(a\,|\,s)-\pi_{k}(a\,|\,s))=V_{\mu}^{\pi}-V_{\mu}^{\pi_{k}}.$
@@
Since $V_{\mu}^{\pi_{k}}$ are non-decreasing by Lemma 6.1 and KL-divergence is non-negative, we conclude that
@@
For given $\eta,\epsilon>0$, there exist $\pi$ such that $V^{\star}_{+,\mu}-V^{\pi}<\epsilon/2$ since $V^{\star}_{+,\mu}<\infty$. Then, by previous inequality, there exist $\pi_{k}$ such that $V^{\pi}-V^{\pi_{k}}_{\mu}<\epsilon/2$. Thus, we have $V^{\star}_{+,\mu}-V^{\pi_{k}}_{\mu}<\epsilon$. Since this holds for arbitrary $\epsilon$, we get $V^{\pi_{k}}_{\mu}\rightarrow V^{\star}_{+,\mu}.$
C.4 Proof of Corollary 6.3
@@
In the previous proof of Theorem 6.2, let $\pi=\pi^{\star}$, and use Corollary A.3 instead of Corollary A.2 and the fact that $V^{\pi}_{\mu}\geq 0$. Then, we obtain the desired result. ∎
C.5 Proof of Theorem 6.4
@@
Lemma C.3.
Under Assumption 2.1, for a given $\mu$ with full support, the natural policy gradient with step size $\eta_{k}>0$ generates a sequence of policies $\{\pi_{k}\}^{\infty}_{k=1}$ satisfying,
@@
Proof of Lemma C.3.
In the previous proof of Theorem 6.2, we showed that
$\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(s,a)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))+\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(s,a)(\pi(a\,|\,s)-\pi_{k}(a\,|\,s))$
@@
$\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(s,a)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))$
$=\sum_{s\in\mathcal{S}}\frac{\delta_{\mu}^{\pi}}{\delta_{\mu}^{\pi_{k+1}}}\,\delta_{\mu}^{\pi_{k+1}}\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(s,a)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))$
$\geq\left\lVert\frac{\delta_{\mu}^{\pi}}{\delta_{\mu}^{\pi_{k+1}}}\right\rVert_{\infty}\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi_{k+1}}\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(s,a)(\pi_{k}(a\,|\,s)-\pi_{k+1}(a\,|\,s))$
@@
where the first inequality is from Lemma 6.1, and the last equality comes from Corollary A.2.
For the second term, by Corollary A.2,
$\sum_{s\in\mathcal{S}}\delta_{\mu}^{\pi}(s)\sum_{a\in\mathcal{A}}Q^{\pi_{k}}(s,a)(\pi(a\,|\,s)-\pi_{k}(a\,|\,s))=V_{\mu}^{\pi}-V_{\mu}^{\pi_{k}}.$
we obtain the desired result after substitution.
We are now ready to prove Theorem 6.4
@@
Since $\delta_{\mu}^{\pi_{k}}(s)\geq\mu(s)$ for all $s\in\mathcal{S}$, $\vartheta^{\pi}_{k}\leq\vartheta^{\pi}_{\mu}$ for all $k\geq 0$. $U^{\pi}_{k+1}-U^{\pi}_{k}\leq 0$ for all $k\geq 0$ by Lemma C.2, and by Lemma C.3,